OpenStudy (anonymous):

A 25000 Kg airplane takes off from rest and reaches 5000m altitude and cruises at 900km/h in 8 minutes .Find Avg Power ?

OpenStudy (anonymous):

My Work: Pavg = W/t W = Delta K.E W = 25000 * 250*250 / 2 t = 8 * 60

OpenStudy (anonymous):

i got Pavg = 1.6 MW but the answer shuld be 4.2 MW

terenzreignz (terenzreignz):

Hang on a bit...

OpenStudy (vincent-lyon.fr):

You have forgotten the gain in gravitational potential energy.

OpenStudy (anonymous):

We have Consider P.E here ?

OpenStudy (vincent-lyon.fr):

Of course! How could you raise the altitude of the plane without providing energy?!

OpenStudy (anonymous):

Yup...i got..it...thx..)

OpenStudy (anonymous):

but @Vincent-Lyon.Fr we knw change in K.E is Work Done..Why cant we Use it here..)

OpenStudy (anonymous):

Is it then mgh+delta K.E /t ?

OpenStudy (vincent-lyon.fr):

You can, but there are two works to take into account: - motive work by the engine (positive) - negative work by the weight so change in KE = Wengine - Mgh or W engine = change in KE + change in PE

OpenStudy (anonymous):

Ok...Thxx....

OpenStudy (vincent-lyon.fr):

yw :)

OpenStudy (vincent-lyon.fr):

I see you are now getting to the energy concept. Soon you will be able to solve problems of mechanics by this method.